Due to erratic sales of its sole product—a high-capacity battery for laptop computers—PEM, Inc., has been experiencing financial difficulty for some time. The company’s contribution format income statement for the most recent month is given below:

Sales (13,100 units × $30 per unit) $ 393,000
Variable expenses 196,500
Contribution margin 196,500
Fixed expenses 219,000
Net operating loss $ (22,500 )

Required:

1. Compute the company’s CM ratio and its break-even point in unit sales and dollar sales.

2. The president believes that a $6,800 increase in the monthly advertising budget, combined with an intensified effort by the sales staff, will result in an $84,000 increase in monthly sales. If the president is right, what will be the increase (decrease) in the company’s monthly net operating income?

3. Refer to the original data. The sales manager is convinced that a 10% reduction in the selling price, combined with an increase of $38,000 in the monthly advertising budget, will double unit sales. If the sales manager is right, what will be the revised net operating income (loss)?

4. Refer to the original data. The Marketing Department thinks that a fancy new package for the laptop computer battery would grow sales. The new package would increase packaging costs by 0.60 cents per unit. Assuming no other changes, how many units would have to be sold each month to attain a target profit of $4,300?

5. Refer to the original data. By automating, the company could reduce variable expenses by $3 per unit. However, fixed expenses would increase by $58,000 each month.

a. Compute the new CM ratio and the new break-even point in unit sales and dollar sales.

b. Assume that the company expects to sell 20,800 units next month. Prepare two contribution format income statements, one assuming that operations are not automated and one assuming that they are. (Show data on a per unit and percentage basis, as well as in total, for each alternative.)

c. Would you recommend that the company automate its operations (Assuming that the company expects to sell 20,800)?
Answer #1

1.CM Ratio = Contribution Margin/Sales

= 196,500/393,000

= 50%

Break even point in unit sales = Fixed costs/Contribution Margin per unit

= 219,000/15

= 14,600 units

Dollar sales = 219,000/50%

= $438,000

2.Increase in Net Operating Income = Increase in Contribution Margin – Increase in cost

= 84,000*50% - 6,800

= $35,200

3.Net operating Income = 26,200*(27-15) – 219,000-38,000

= $57,400

4.Target Profit = $4,300

Fixed costs = 219,000

Target Contribution Margin = $223,300

Units required to be sold = 223,300/(30-15-0.6)

= 15,506.94 units

5.CM Ratio = (30-12)/30 = 60%

Break even point in unit sales = (219,000+58,000)/18 = 15,388.89 units

Dollar sales = 277,000/60%

= $461,666.67

b.

Without Automation

With Automation

Per unit

Total

%

Per unit

Total

%

Sales

30

624,000

100.00%

30

624,000

100.00%

Variable expenses

15

312,000

50.00%

12

249,600

40.00%

Contribution Margin

15

312,000

50.00%

18

374,400

60.00%

Fixed Expenses

10.52885

219,000

35.10%

13.31731

277,000

44.39%

Net operating income

4.471154

93,000

14.90%

4.682692

97,400

15.61%

Yes, should automate